OpenStudy (anonymous):

How to solve quadratic equations... specifically this one: h = -16t2 + 36t + 4.

OpenStudy (here_to_help15):

I'll help :)

OpenStudy (zale101):

use the quadratic formula :)

OpenStudy (anonymous):

Ok i understand that... I know the quadratic formula.. im asking if someone can help me step by step, Otherwise I wouldn't be asking now would i ? @Zale101

OpenStudy (here_to_help15):

:)

OpenStudy (anonymous):

Ok so are you guys just going to type smiley faces or will someone genuinely help me?

OpenStudy (here_to_help15):

i will help just dont gget mad at me ok?

OpenStudy (zale101):

 ax^2+bx+c=0 has important variables a, b, c that can be plugged in the quadratic formula. Your quadratic equation is h = -16t2 + 36t + 4. -16t^2 + 36t + 4=0 it's in the form of ax^2+bx+c=0 but a is -16 b is 36 and c is 4 plug in to the quadratic formula, which you are already familiar with.